The distribution of times that a companys service technicians take to respond to trouble calls is normal with mean μ and standard deviation 0.25 hours. The company advertises that its service technicians take an average of 2 hours to respond to trouble calls from customers. We suspect that the average time is really more than 2 hours. From a random sample of 25 trouble calls, the average time service technicians took to respond was 2.10 hours. How strong is the evidence against the companys claim? Based on these data, the P -value of the appropriate test is (express answer with 4 places after decimal) Answer:
Answer #1

Solution:

Given:

Mean =\mu = 2

Standard Deviation = \sigma= 0.25

Sample size = n = 25

Sample mean = \bar{x}= 2.10

We have to test if the average service time is more than 2 hours.

Thus hypothesis are:
H_{0}: \mu = 2

Vs

H_{1}: \mu > 2

We have to find p-value:

p-value = P( Z> z test statistic )

where

o/n

z=\frac{ 2.10-2 }{0.25 /\sqrt{25}}

z=\frac{ 0.10 }{0.25 / 5}

z=\frac{ 0.10 }{0.05 }

z= 2.00

Thus

p-value = P( Z> 2.00)

p-value = 1 - P( Z < 2.00)

Look in z table for z = 2.0 and 0.00 and find corresponding area.

2 .00 .01 10.0 .50 po .5040 10.1 5398 5438 10.2 .5723 5832 10.3 6179 6217 .64 .6591 6915 .6950 77 7291 1.7480 7611 7881 7910

P( Z< 2.00) = 0.9772

thus

p-value = 1 - P( Z < 2.00)

p-value = 1 - 0.9772

p-value = 0.0228